OpenStudy (anonymous):

Please help: The sum of twice a number and 4 less than the number is the same as the difference between -32 and the number. What is the number?

OpenStudy (anonymous):

Try to rewrite the question, replacing 'the number' with x; 'sum' with +, 'the same as' with =.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

is it x = -32 - 4 x = -36?

OpenStudy (anonymous):

If we look at the first part, it says 'the sum of twice a number, and 4 less than the number'. How would you write 'twice a number'?

OpenStudy (anonymous):

n*2?

OpenStudy (anonymous):

or 2n

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

Yep! What about '4 less than the number'?

OpenStudy (anonymous):

2n - 4?

myininaya (myininaya):

sum means add twice a number means 2*n where n is a number 4 less than a number means n-4 so the sum of those is 2n+(n-4) the same as means equal difference of -32 and n is -32-n 2n+(n-4)=-32-n

OpenStudy (anonymous):

Oh, okay. Thank you